The speed of the freight train is 23 mph slower than that of a passenger train. The freight train travels 250 miles in the same time it takes the passenger train to travel 450 miles. Find the speed of the passenger train.
Answer by mananth(16946)

freight train speed = x mph
passenger train speed = x+23 mph
..
distance by freight train = 250 miles
distrance by passenger = 450 miles.
..
Time taken by both = same
..
250/x = 450/(x+23)
250(x+23)=450x
250x+5750=450x
450x-250x=5750
200x = 5750
/200
x= 5750/200
x=28.75 mph freight train
passenger train speed = 28.75+23=51.75mph